OpenStudy (anonymous):

−4(x + 1) − 3 = −3(x − 4) is it 0?

OpenStudy (anonymous):

no please show your work

OpenStudy (anonymous):

Mkay so first I do the parenthese right ? which i get 5 .-. i think lol

OpenStudy (anonymous):

how? you cant add the variable to the constant you want to first distribute the numbers distribution property of multiplication a(b+c) = ab + ac

jimthompson5910 (jim_thompson5910):

I think somehow you saw the 4 outside the parenthesis and added it with the 1 inside the parenthesis to get 4+1 = 5

jimthompson5910 (jim_thompson5910):

not 100% sure though

OpenStudy (anonymous):

mkay so what do i distribute

jimthompson5910 (jim_thompson5910):

−4(x + 1) would turn into -4 times x + (-4) times 1 and that turns into -4x + (-4) or -4x - 4

jimthompson5910 (jim_thompson5910):

so to sum things up −4(x + 1) becomes -4x - 4

jimthompson5910 (jim_thompson5910):

the rule of distribution is that you multiply the outer number by everything inside, then you add up those products

OpenStudy (anonymous):

okay imma do that

jimthompson5910 (jim_thompson5910):

distribute out −3(x − 4) and tell me what you get

OpenStudy (anonymous):

it has to be -19

jimthompson5910 (jim_thompson5910):

what do you get when you distribute −3(x − 4)

jimthompson5910 (jim_thompson5910):

-19 is the correct final answer, hopefully you see how/why